# Understanding Flow Behaviors of Supercooled Liquids by Embodying Solid-Liquid Duality at Particle Level

### v1

@@ IV.2 Structural basis for solid-liquid duality
Figure 19: Spatial distributions of the hardness $H$ at various $\dot{\gamma}$ for both samples. (a) – (d) show the configuration slices selected at different $\dot{\gamma}$ for sample A. (e) – (h) show the configuration slices selected at different $\dot{\gamma}$ for sample B. In all panels, the color of each particle denotes its value of $H$ with a common color bar. The boundaries between liquid-like (fast) and solid-like (slow) groups are denoted by black dashed lines. A nice agreement between the boundaries and the particles with $H$ close to 00 is seen in all panels.
In the preceding subsection, $\tau_{\mathrm{LC}}$ extracted from different $\dot{\gamma}$ have been quantified through a unified viewpoint of structure. Moreover, in Sec. II, we reveal a firm relation between the local response and the local dynamics. Therefore, one may expect a structural basis for the local response. Particularly, we have demonstrated that the local response has two forms: solid-like and liquid-like. As $\dot{\gamma}$ increases, the ratio of these two components varies, corresponding to the flattening of the form of $\tau_{\mathrm{sh}}(\tau_{\mathrm{LC}})$. Here, we aim to comprehend such variation with the assistance of structural parameters. The key is to find a structural criterion, which is universal for all shear rates, for classifying particles into solid-like and liquid-like groups.
@@
Here, we still use the feature vector introduced in Eq. 13 to characterize the local structure of each particle. To construct the data set for the SVM model, we select $10$ configurations from each $\dot{\gamma}$, ranging from the equilibrium state to the shear-thinning regime. For each configuration, particles are classified as slow and fast using the method given in Sec. II. Slow particles are assigned to the solid-like group $G_{\mathrm{s}}$, while fast particles are assigned to the liquid-like group $G_{\mathrm{l}}$. The number of elements in $G_{\mathrm{s}}$ and that in $G_{\mathrm{l}}$ are set to be equal. The data set is randomly partitioned into a training set and a test set with a ratio of $8:2$. Particles in $G_{\mathrm{s}}$ are labeled as $5$, while those in $G_{\mathrm{l}}$ are labeled as $-5$. Then, a SVM algorithm with a linear kernel is employed on the training set. The penalty factor is finely tuned to achieve the optimal classification. With this operation, we construct a hyperplane in the high dimensional space defined by the feature vector to separate the two categories of particles. We examine the accuracy of the classification with the test set, and the result is $82\%$.
@@
To validate our SVM model, it is crucial to check if it can correctly capture the solid-like and liquid-like regions. In Fig. 19, we plot several slices of both systems at different $\dot{\gamma}$. The upper and lower panels represent sample A and sample B, respectively. Particles are colored based on their values of $H$. We also plot the boundary between fast and slow regions for all slices according to the method given in Sec. II with black dashed lines. It is seen that the two methods give highly consistent spatial features of the solid-liquid duality in local response.
@@
There is another important observation from Fig. 19. Firstly, we recall that the model for determining $H$ is universal and valid for all shear rates. Then, as shown in Fig. 19, we find that more particles migrate to the $H>0$ side with the growth of $\dot{\gamma}$. To quantify this observation, we calculate the average value of $H$ over all particles $\bar{H}$ for each $\dot{\gamma}$ and both samples. The results are given in Fig. 20 (a) and (b). As $\dot{\gamma}$ increases, $\bar{H}$ starts to increase from a negative value, passes through zero, and finally becomes positive. It is interesting to see that the shear rate at which $\bar{H}$ crosses zero just coincides with the crossover regime. We can evaluate the percentage of slow particles by calculating the proportion of particles with $H>0$. The results, denoted as $p_{H}$, are shown in Fig. 20 (c) and (d) for sample A and sample B, respectively. Meanwhile, we replot the results of $p_{\mathrm{slow}}$ shown in Fig. 13. It is seen that $p_{H}$ and $p_{\mathrm{slow}}$ are consistent with each other. This agreement provides a structural view for the shear-induced changes in dynamics and local response.
@@
Figure 20: (a) and (b) show the variation of the configuration-averaged $H$, denoted as $\bar{H}$, with respect to $\dot{\gamma}$ for sample A and sample B, respectively. In both (a) and (b), vertical yellow bands denote the crossover regime. (c) and (d) show the proportion $p_{H}$ of particles with $H>0$ (circles and squares) for sample A and sample B, respectively. Squares denote the results at the equilibrium state, and circles denote the results at other shear rates. In both (c) and (d), the results of $p_{\mathrm{slow}}$ (diamonds), which have been shown in Fig. 13, are replotted.
@@
The result discussed in this subsection makes an advance to the preceding subsection. In the preceding subsection, we find the correlation between structure and $\tau_{\mathrm{LC}}$. This correlation is somewhat natural: By definition, $\tau_{\mathrm{LC}}$ is calculated by using the ICE method and stopping external shear. In this way, the convection and thermal noise are removed, which highlights the role of the transient configuration in determining $\tau_{\mathrm{LC}}$. This correlation is irrelevant to local response to shear and the form of $\tau_{\mathrm{sh}}(\tau_{\mathrm{LC}})$. In this subsection, by utilizing the same feature vector, we build the connection between local response and local structure. Note that the form of local response involves the interplay among external shear, thermal activation and structure. The effectiveness of $H$ in classification and the variations of $\bar{H}$ and $p_{H}$ with $\dot{\gamma}$ further emphasize the importance of local structure in determining the rheology of supercooled liquids. Summarizing all results in this section, we establish the structural basis for the dynamical and rheological behaviors discussed in Secs. II and III.
V DISCUSSION: A PEL PERSPECTIVE
@@
In Secs. IIIV, we find the connection from the local configuration to the local dynamics and response, and finally to the flow behaviors. Following this line of thought, one may expect to formulate the behaviors of $\tau_{\mathrm{LC}}$, the form of $\tau_{\mathrm{sh}}(\tau_{\mathrm{LC}})$, and the duality in local response, which compose the foundation for understanding the flow of supercooled liquids in our framework, with theoretical tools involving local configuration and dynamics. The key issue is to elucidate how the structure, flow and thermal effect collectively affect the local dynamics. In this section, this problem will be discussed by utilizing and modestly extending several critical ingredients of the PEL picture.
V.1 PEL perspective on $\tau_{\mathrm{LC}}$ and $\tau_{\mathrm{sh}}$
Supercooled liquids are featured by amorphous configuration and slow dynamics [15, 56, 80]. PEL provides a useful way to bridge these two features [47, 48, 49]. In this picture, the potential energy of the system, which is a function of coordinates of all particles, defines a complex $3N$-dimensional surface composed of numerous basins. For a supercooled liquid, the system is trapped in basins for most of the time, and undergoes inter-basin jumps occasionally by thermal activation [47, 58, 49].
Despite the convenience and clearness in concept, the effectiveness and complexity of PEL method get worse for big systems [47], because the inter-basin jump of big systems is very frequent and its effect is not so sharp. On the other hand, trap model provides an effective approach for understanding the slow dynamics [109, 132]. In this model, the whole system is divided into many coherent subsystems. For each subsystem, the picture of PEL is applied: the subsystem is confined in basins (traps) separated by barriers that can be overcome by thermal activation. The dynamics of the whole system is related to the statistics of the configuration of subsystem.
PEL is tilted as shear is applied. As strain increases, the degree of tilt and elastic stress are both enhanced, until mechanical instability emerges [50, 133]. Then, the shear stress is released, and the system goes into next minimum. For glasses under steady shear, such stress buildup and strain-activated process take place successively, which plays a key role in the non-Newtonian behavior [134].
@@ V.1 PEL perspective on $\tau_{\mathrm{LC}}$ and $\tau_{\mathrm{sh}}$
Figure 21: (a) A review of the flow conditions for measuring $\tau_{\mathrm{LC}}$ and $\tau_{\mathrm{sh}}$ under steady shear with respect to a configuration at time $t_{1}$. (b) Illustration of the evolution of PEL for measuring $\tau_{\mathrm{sh}}$ under steady shear. Here, the PEL shows an increasing tilt as time goes, represented by the change in shape from dashed lines to solid lines. (c) Illustration of the evolution of PEL for measuring $\tau_{\mathrm{LC}}$. Here, the tilting of PEL stops at $t=t_{1}$ due to the abrupt cession of flow.
@@
In this subsection, we will model the behaviors of $\tau_{\mathrm{LC}}$ and $\tau_{\mathrm{sh}}$ by combining the ideas discussed in the above three paragraphs. Considering that both $\tau_{\mathrm{LC}}$ and $\tau_{\mathrm{sh}}$ are defined as the time for a reference particle to jump out, we can assume a basin for the particle to represent the potential energy barrier it experiences [26]. The shape of the basin is determined by the local configuration of the reference particle. When shear is applied, the basin is tilted. In Fig. 21 (a), we review the flow conditions for extracting $\tau_{\mathrm{LC}}$ and $\tau_{\mathrm{sh}}$ with respect to the configuration at $t=t_{1}$. Figure 21 (b) illustrates the evolution of PEL for measuring $\tau_{\mathrm{sh}}$. In this case, the shear is steady, and the PEL tilts more and more with time before the mechanical instability. Figure 21 (c) illustrates the evolution of PEL for measuring $\tau_{\mathrm{LC}}$. Here, the external shear stops at $t=t_{1}$. Consequently, the basin also stops tilting at $t=t_{1}$. Figure 21 elucidates the difference between $\tau_{\mathrm{LC}}$ and $\tau_{\mathrm{sh}}$. $\tau_{\mathrm{sh}}$ describes the local relaxation in flow. It is determined by the combined effect of thermal activation and convection. For $\tau_{\mathrm{LC}}$, the convection is eliminated by stopping external shear and, thus, it is determined by the thermal activation from the distorted configuration at $t=t_{1}$. From the above discussion, it is seen that $\tau_{\mathrm{LC}}$, rather than $\tau_{\mathrm{sh}}$, directly reflects the local configuration and the particle-level energy barrier.
@@
where $\tau_{\mathrm{LC,eq}}$ denotes the $\tau_{\mathrm{LC}}$ at equilibrium. With Eq. 14 and the spectrum of $\tau_{\mathrm{LC,eq}}$, the distribution of $E_{\mathrm{b}}$ can be obtained.
In the presence of external shear, the local structure undergoes deformation at a strain rate of $\dot{\gamma}$. During the deformation, the local structure will store potential energy until the activation takes place. This deformation energy is denoted as $E(\gamma)$, where $\gamma$ is the local strain. Taking into account both thermal activation and shear-induced deformation, the probability for a particle to escape the basin within $\mathrm{d}t$ is given by:
@@
Here, the form of $E(\gamma)$ is important. For small strain, $E(\gamma)$ is harmonic. However, as $\gamma$ increases, $E(\gamma)$ gradually deviates from the harmonic form. As predicted by the catastrophe theory, when $\gamma$ approaches the edge of basin $\gamma_{\mathrm{b}}$, the energy increment $\Delta E$ scales with the residual load $\gamma_{\mathrm{b}}-\gamma$ as $\Delta E\propto(\gamma_{\mathrm{b}}-\gamma)^{3/2}$ [137]. Moreover, Maloney and Lacks demonstrate that this scaling remains valid well beyond the vanishing regime of $\gamma_{\mathrm{b}}-\gamma$, even at finite temperatures [138]. Considering the behaviors of $E(\gamma)$ at small and large strains, we model its form as:
[…]
@@
where $G$ is the local shear modulus, $\Omega$ is the activation volume, $\gamma_{\mathrm{x}}$ is the crossover strain, and $A$ is a numerical factor. Figure 22 illustrates the form of $E(\gamma)$. By requiring that both $E(\gamma)$ and its first-order derivative are continuous at $\gamma=\gamma_{\mathrm{x}}$, the three parameters, $\gamma_{\mathrm{x}}$, $\gamma_{\mathrm{b}}$ and $A$, can be reduced to a single parameter. As illustrated in Fig. 22, we can denote the value of $E(\gamma_{\mathrm{x}})$ by $\alpha E_{\mathrm{b}}$, where $\alpha\,\left(\alpha\in(0,1)\right)$ is employed as the free adjustable parameter that represents $\gamma_{\mathrm{x}}$, $\gamma_{\mathrm{b}}$ and $A$.
@@
Figure 22: Illustration of the deformation energy $E(\gamma)$ given by Eq. 16. The harmonic part is colored by light red, and the catastrophe part is colored by yellow. $\gamma_{\mathrm{x}}$ denotes the boundary between these two parts. $\alpha E_{\mathrm{b}}$ ($10<\alpha<$ 1) gives the value of $E(\gamma=\gamma_{\mathrm{x}})$.
@@
Besides the free parameter $\alpha$, other parameters in Eq. 15 are set as follows: $\Gamma^{-1}$ is determined from the MSD data at equilibrium as the time at which the particle enters the plateau regime. $G$ is set to the effective modulus $G_{\mathrm{eff}}$ defined in Sec. III. $\Omega$ represents the size that will be involved for the reference particle to jump out. At the current stage, we estimate $\Omega$ by the volume of the cage, i.e., the sphere whose radius is given by the distance of the first minimum of $g(r)$. For the sake of simplicity, we assume that all these parameters do not vary with $\dot{\gamma}$. Note that our assumption on $\Omega$ is oversimplified. As demonstrated in Sec. III, the component of solid-like parts increases with $\dot{\gamma}$, which may result in the enhancement of the correlation between particles. We will revisit this point in Sec. V.3.
@@
Figure 23: $\tau_{\mathrm{sh}}(\tau_{\mathrm{LC}})$ calculated from the PEL model (lines) for sample A (a) and sample B (b) at various $\dot{\gamma}$. The results obtained from MD data (symbols) are also plotted. In each panel, the results corresponding to the same $\dot{\gamma}$ are displayed in the same color.
@@
Assuming that the basin starts deforming at $t=0$, the probability that a particle jumps at $t$ can be derived with Eq. 15 and is expressed as:
@@
×$\times\exp\left(-\Gamma\int_{0}^{t}\exp\left[-\frac{E_{\mathrm{b}}-E(\dot{\gamma}t^{\prime})}{k_{\mathrm{B}}T}\right]\mathrm{d}t^{\prime}\right).$
@@
Setting $\dot{\gamma}=0$, Eq. 18 reduces to Eq. 14 and gives the value of $\tau_{\mathrm{LC}}$ at equilibrium. For $\dot{\gamma}>0$, $\bar{t}(E_{\mathrm{b}},\dot{\gamma})$ gives $\tau_{\mathrm{sh}}$ according to the picture shown in Fig. 21. Therefore, with Eq. 18, we can get the form of $\tau_{\mathrm{sh}}(\tau_{\mathrm{LC}})$. Figure 23 shows $\tau_{\mathrm{sh}}(\tau_{\mathrm{LC}})$ found by the PEL model at various $\dot{\gamma}$ for both samples. For comparison, we also plot the MD results of $\tau_{\mathrm{sh}}(\tau_{\mathrm{LC}})$. It is seen that the PEL model captures the general characteristics of $\tau_{\mathrm{sh}}(\tau_{\mathrm{LC}})$. Particularly, its form exhibits a greater flattening with $\dot{\gamma}$. From the viewpoint of this model, this flattening is due to the increasingly dominant role of shear deformation compared to thermal activation as $\dot{\gamma}$ increases.

### v2

@@ IV.2 Structural basis for solid-liquid duality
In the preceding subsection, $\tau_{\mathrm{LC}}$ extracted from different $\dot{\gamma}$ have been quantified through a unified viewpoint of structure. Moreover, in section II, we reveal a firm relation between the local response and the local dynamics. Therefore, one may expect a structural basis for the local response. Particularly, we have demonstrated that the local response has two forms: solid-like and liquid-like. As $\dot{\gamma}$ increases, the ratio of these two components varies, corresponding to the flattening of the form of $\tau_{\mathrm{sh}}(\tau_{\mathrm{LC}})$. Here, we aim to comprehend such variation with the assistance of structural parameters. The key is to find a structural criterion, which is universal for all shear rates, for classifying particles into solid-like and liquid-like groups.
@@
Here, we still use the feature vector introduced in equation 13 to characterize the local structure of each particle. To construct the data set for the SVM model, we select $10$ configurations from each $\dot{\gamma}$, ranging from the equilibrium state to the shear-thinning regime. For each configuration, particles are classified as slow and fast using the method given in section II. Slow particles are assigned to the solid-like group $G_{\mathrm{s}}$, while fast particles are assigned to the liquid-like group $G_{\mathrm{l}}$. The number of elements in $G_{\mathrm{s}}$ and that in $G_{\mathrm{l}}$ are set to be equal. The data set is randomly partitioned into a training set and a test set with a ratio of $8:2$. Particles in $G_{\mathrm{s}}$ are labeled as $5$, while those in $G_{\mathrm{l}}$ are labeled as 5-5. Then, a SVM algorithm with a linear kernel is employed on the training set. The penalty factor is finely tuned to achieve the optimal classification. With this operation, we construct a hyperplane in the high dimensional space defined by the feature vector to separate the two categories of particles. We examine the accuracy of the classification with the test set, and the result is $82\%$.
@@
To validate our SVM model, it is crucial to check if it can correctly capture the solid-like and liquid-like regions. In figure 19, we plot several slices of both systems at different $\dot{\gamma}$. The upper and lower panels represent sample A and sample B, respectively. Particles are colored based on their values of $H$. We also plot the boundary between fast and slow regions for all slices according to the method given in section II with black dashed lines. It is seen that the two methods give highly consistent spatial features of the solid-liquid duality in local response.
@@
There is another important observation from figure 19. Firstly, we recall that the model for determining $H$ is universal and valid for all shear rates. Then, as shown in figure 19, we find that more particles migrate to the $H>0$ side with the growth of $\dot{\gamma}$. To quantify this observation, we calculate the average value of $H$ over all particles $\bar{H}$ for each $\dot{\gamma}$ and both samples. The results are given in figure 20 (a) and (b). As $\dot{\gamma}$ increases, $\bar{H}$ starts to increase from a negative value, passes through zero, and finally becomes positive. It is interesting to see that the shear rate at which $\bar{H}$ crosses zero just coincides with the crossover regime. We can evaluate the percentage of slow particles by calculating the proportion of particles with $H>0$. The results, denoted as $p_{H}$, are shown in figure 20 (c) and (d) for sample A and sample B, respectively. Meanwhile, we replot the results of $p_{\mathrm{slow}}$ shown in figure 13. It is seen that $p_{H}$ and $p_{\mathrm{slow}}$ are consistent with each other. This agreement provides a structural view for the shear-induced changes in dynamics and local response.
@@
Figure 20: (a) and (b) show the variation of the configuration-averaged $H$, denoted as $\bar{H}$, with respect to $\dot{\gamma}$ for sample A and sample B, respectively. In both (a) and (b), vertical yellow bands denote the crossover regime. (c) and (d) show the proportion $p_{H}$ of particles with $H>0$ (circles and squares) for sample A and sample B, respectively. Squares denote the results at the equilibrium state, and circles denote the results at other shear rates. In both (c) and (d), the results of $p_{\mathrm{slow}}$ (diamonds), which have been shown in figure 13, are replotted.
@@
The result discussed in this subsection makes an advance to the preceding subsection. In the preceding subsection, we find the correlation between structure and $\tau_{\mathrm{LC}}$. This correlation is somewhat natural: By definition, $\tau_{\mathrm{LC}}$ is calculated by using the ICE method and stopping external shear. In this way, the convection and thermal noise are removed, which highlights the role of the transient configuration in determining $\tau_{\mathrm{LC}}$. This correlation is irrelevant to local response to shear and the form of $\tau_{\mathrm{sh}}(\tau_{\mathrm{LC}})$. In this subsection, by utilizing the same feature vector, we build the connection between local response and local structure. Note that the form of local response involves the interplay among external shear, thermal activation and structure. The effectiveness of $H$ in classification and the variations of $\bar{H}$ and $p_{H}$ with $\dot{\gamma}$ further emphasize the importance of local structure in determining the rheology of supercooled liquids. Summarizing all results in this section, we establish the structural basis for the dynamical and rheological behaviors discussed in sections II and III.
V DISCUSSION: A MODEL ANALYSIS
@@
In sections IIIV, we find the connection from the local configuration to the local dynamics and response, and finally to the flow behaviors. Following this line of thought, one may expect to formulate the behaviors of $\tau_{\mathrm{LC}}$, the form of $\tau_{\mathrm{sh}}(\tau_{\mathrm{LC}})$, and the duality in local response, which compose the foundation for understanding the flow of supercooled liquids in our framework, with theoretical tools involving local configuration and dynamics. The key issue is to elucidate how the structure, flow and thermal effect collectively affect the local dynamics. In this section, this problem will be discussed with a simple model employing the picture of shear-facilitated activation at the semiquantitative level.
V.1 Shear-facilitated-activation picture of $\tau_{\mathrm{LC}}$ and $\tau_{\mathrm{sh}}$
Studies of dynamical processes in supercooled liquids typically conceptualize the structure as a basin (or trap) that can be overcome by thermal activation to realize structural relaxation [58, 129, 130]. Such an activation picture has been extensively applied at the local level, resulting in the trap model [49, 131], and even at the particle level, leading to a series of studies by Liu et al. wherein a “softness” parameter is introduced [28, 132]. As previously noted, particles in supercooled liquids are confined within cages for the majority of time, exhibiting intermittent hoppings to new environments. Thus, an energy basin that temporarily confines an individual particle can be unambiguously defined, whereby the rearrangement corresponds to the thermal activations out of the basin.
In the presence of external shear, the configuration is distorted, implying that the energy basin is tilted [24], which facilitates the out-of-basin motion on average [133, 134]. An intriguing interplay between thermal activation and shear deformation emerges under such conditions. The incorporation of both effects was pioneered in the soft glassy rheology (SGR) model [23] and certain EPMs [113], where mesoscopic regions are considered. Since such shear-induced distortion also manifests at the cage level, the interplay between thermal activation and shear deformation should be applicable at this scale.
@@ V.1 Shear-facilitated-activation picture of $\tau_{\mathrm{LC}}$ and $\tau_{\mathrm{sh}}$
Figure 21: (a) A review of the flow conditions for measuring $\tau_{\mathrm{LC}}$ and $\tau_{\mathrm{sh}}$ under steady shear with respect to a configuration at time $t_{1}$. (b) Illustration of the evolution of the basin for measuring $\tau_{\mathrm{sh}}$ under steady shear. Here, the basin shows an increasing tilt as time goes, represented by the change in shape from dashed lines to solid lines. (c) Illustration of the evolution of the basin for measuring $\tau_{\mathrm{LC}}$. Here, the tilting of basin stops at $t=t_{1}$ due to the abrupt cession of flow.
@@
In this subsection, we will model the behaviors of $\tau_{\mathrm{LC}}$ and $\tau_{\mathrm{sh}}$ by combining the ideas discussed in the above two paragraphs. Considering that both $\tau_{\mathrm{LC}}$ and $\tau_{\mathrm{sh}}$ are defined as the time for a reference particle to jump out, we can assume a basin for the particle to represent the potential energy barrier it experiences [28]. The shape of the basin is determined by the local configuration of the reference particle. When shear is applied, the basin is tilted. In figure 21 (a), we review the flow conditions for extracting $\tau_{\mathrm{LC}}$ and $\tau_{\mathrm{sh}}$ with respect to the configuration at $t=t_{1}$. Figure 21 (b) illustrates the evolution of basin for measuring $\tau_{\mathrm{sh}}$. In this case, the shear is steady, and the basin tilts more and more with time before the activation or the mechanical instability occurs. Figure 21 (c) illustrates the evolution of basin for measuring $\tau_{\mathrm{LC}}$. Here, the external shear stops at $t=t_{1}$. Consequently, the basin also stops tilting at $t=t_{1}$. Figure 21 elucidates the difference between $\tau_{\mathrm{LC}}$ and $\tau_{\mathrm{sh}}$. $\tau_{\mathrm{sh}}$ describes the local relaxation in flow. It is determined by the combined effect of thermal activation and convection. For $\tau_{\mathrm{LC}}$, the convection is eliminated by stopping external shear and, thus, it is determined by the thermal activation from the distorted configuration at $t=t_{1}$. From the above discussion, it is seen that $\tau_{\mathrm{LC}}$, rather than $\tau_{\mathrm{sh}}$, directly reflects the local configuration and the particle-level energy barrier.
@@
where $\tau_{\mathrm{LC,eq}}$ denotes the $\tau_{\mathrm{LC}}$ at equilibrium. With equation 14 and the distribution of $\tau_{\mathrm{LC,eq}}$, the distribution of $E_{\mathrm{b}}$ can be obtained. As evident from the spatial distribution of $\tau_{\mathrm{LC}}$ (figure 2), while energy barriers $E_{\mathrm{b}}$ are defined for individual particles, they exhibit a smooth spatial variation with clustering features.
In the presence of external shear, the local structure undergoes deformation at a strain rate of $\dot{\gamma}$. During the deformation, the local structure will store potential energy until the activation takes place. This deformation energy is denoted as $E(\gamma)$, where $\gamma$ is the local strain. Taking into account both thermal activation and shear effect, the probability for a particle to escape the basin within $\mathrm{d}t$ is given by:
@@
Here, the form of $E(\gamma)$ is important. $E(\gamma)$ is harmonic at small strains. As $\gamma$ increases, $E(\gamma)$ gradually deviates from the harmonic form. As predicted by the catastrophe theory, when $\gamma$ approaches the edge of basin $\gamma_{\mathrm{b}}$, the energy increment $\Delta E$ scales with the residual load $\gamma_{\mathrm{b}}-\gamma$ as $\Delta E\propto(\gamma_{\mathrm{b}}-\gamma)^{3/2}$ [137]. Moreover, Maloney and Lacks demonstrate that this scaling remains valid well beyond the vanishing regime of $\gamma_{\mathrm{b}}-\gamma$, even at finite temperatures [138]. Considering the behaviors of $E(\gamma)$ at small and large strains, we model its form as:
$E(\gamma)=\begin{cases}\frac{1}{2}G\Omega\gamma^{2}&\;\,0\leq\gamma\leq\gamma_{\mathrm{x}}\\ E_{\mathrm{b}}-A(\gamma_{\mathrm{b}}-\gamma)^{3/2}&\gamma_{\mathrm{x}}<\gamma\leq\gamma_{\mathrm{b}}\end{cases},$ (16)
@@
where $G$ is the local shear modulus, $\Omega$ is the activation volume, $\gamma_{\mathrm{x}}$ is the crossover strain, and $A$ is a numerical factor. Figure 22 illustrates the form of $E(\gamma)$. By requiring that both $E(\gamma)$ and its first-order derivative are continuous at $\gamma=\gamma_{\mathrm{x}}$, the three parameters, $\gamma_{\mathrm{x}}$, $\gamma_{\mathrm{b}}$ and $A$, can be reduced to a single parameter. As illustrated in figure 22, we can denote the value of $E(\gamma_{\mathrm{x}})$ by $\alpha E_{\mathrm{b}}$, where $\alpha\,\left(\alpha\in(0,1)\right)$ is employed as the free adjustable parameter that represents $\gamma_{\mathrm{x}}$, $\gamma_{\mathrm{b}}$ and $A$.
@@
Figure 22: Illustration of the deformation energy $E(\gamma)$ given by equation 16. The harmonic part is colored by light red, and the catastrophe part is colored by yellow. $\gamma_{\mathrm{x}}$ denotes the boundary between these two parts. $\alpha E_{\mathrm{b}}$ ($0<\alpha<1$) gives the value of $E(\gamma=\gamma_{\mathrm{x}})$.
@@
Besides the free parameter $\alpha$, other parameters in equation 15 are set as follows: $\Gamma^{-1}$ is determined from the MSD data at equilibrium as the time at which the particle enters the plateau regime. $G$ is set to the effective modulus $G_{\mathrm{eff}}$ defined in section III. $\Omega$ represents the size that will be involved for the reference particle to jump out. At the current stage, we estimate $\Omega$ by the volume of the cage, i.e., the sphere whose radius is given by the distance of the first minimum of $g(r)$. For the sake of simplicity, we assume that all these parameters do not vary with $\dot{\gamma}$. Note that our assumption on $\Omega$ is oversimplified. As demonstrated in section III, the component of solid-like parts increases with $\dot{\gamma}$, which may result in the enhancement of the correlation between jump events. We will revisit this point in section V.3.
@@
Figure 23: $\tau_{\mathrm{sh}}(\tau_{\mathrm{LC}})$ calculated from the model analysis (lines) for sample A (a) and sample B (b) at various $\dot{\gamma}$. The results obtained from MD data (symbols) are also plotted. In each panel, the results corresponding to the same $\dot{\gamma}$ are displayed in the same color.
@@
Assuming that the basin starts deforming at $t=0$, the probability that a particle jumps at $t$ can be derived with equation 15 and is expressed as:
@@
$\times\exp\left(-\Gamma\int_{0}^{t}\exp\left[-\frac{E_{\mathrm{b}}-E(\dot{\gamma}t^{\prime})}{k_{\mathrm{B}}T}\right]\mathrm{d}t^{\prime}\right).$
@@
Setting $\dot{\gamma}=0$, equation 18 reduces to equation 14 and gives the value of $\tau_{\mathrm{LC}}$ at equilibrium. For $\dot{\gamma}>0$, $\bar{t}(E_{\mathrm{b}},\dot{\gamma})$ gives $\tau_{\mathrm{sh}}$ according to the picture shown in figure 21. Therefore, with equation 18, we can calculate the form of $\tau_{\mathrm{sh}}(\tau_{\mathrm{LC}})$. Figure 23 shows $\tau_{\mathrm{sh}}(\tau_{\mathrm{LC}})$ calculated by this approach at various $\dot{\gamma}$ for both samples. For comparison, we also plot the MD results of $\tau_{\mathrm{sh}}(\tau_{\mathrm{LC}})$. It is seen that our model captures the general characteristics of $\tau_{\mathrm{sh}}(\tau_{\mathrm{LC}})$. Particularly, its form exhibits a greater flattening with $\dot{\gamma}$. From the viewpoint of this model, this flattening is due to the increasingly dominant role of shear deformation compared to thermal activation as $\dot{\gamma}$ increases.